The owner of Maumee Ford-Volvo wants to study the relationship between the age of a car and its selling price. Listed below is a random sample of 12 used cars sold at the dealership during the last year.

Car Age (years) Selling Price ($000)
1 11 12.2
2 8 11.0
3 16 4.9
4 18 4.1
5 9 6.7
6 8 13.6
7 10 11.1
8 16 9.0
9 14 9.0
10 18 4.2
11 6 12.1
12 6 10.4

  Click here for the Excel Data File

  1. Determine the regression equation. (Negative values should be indicated by a minus sign. Round your answers to 3 decimal places.)

  1. Estimate the selling price of an 7-year-old car (in $000). (Round your answer to 3 decimal places.)

  1. Interpret the regression equation (in dollars). (Round your answer to the nearest dollar amount.)
Answer #1

Sol:

The regression equation is: y = 14.909 – 0.451 r .

(a).

a = 14.909
b = -0.451


(b).

esimated selling price = 14.909 - (0.451 * 11) = 9.949.

(c) .

For each additional year, the car decreases 451 in value.
(Calculation = 0.451 * 1000 = 451.)

If you Satisfy with Answer, Please give me "Thumb Up". It is very useful for me.
Thank you for your support.